# Planar Collisionless Shock Simulations with Semi-Implicit Particle-in-Cell Model FLEKS

### v1

@@ 2.1 Local 1D Simulations
Three locations, centered downstream at $x=-2.5\,d_{i,\text{up}}$ (Figure 1i-l), across the shock at $x=2.5\,d_{i,\text{up}}$ (Figure 1m-p), and upstream in front of the shock at $x=7.5\,d_{i,\text{up}}$ (Figure 1q-t), each spanning $5\,d_{i,\text{up}}$, are selected to examine the velocity space distributions. The three-dimensional velocity space distributions are projected onto the $V_{x}-V_{y}$ and $V_{x}-V_{z}$ planes, where the z-axis is approximately aligned with the background magnetic field.
It is important to recognize that this 1D simulation neglects parallel heating along the z-direction. Consequently, no proton thermalization occurs along $V_{iz}$ in the downstream and ramp regions (Figure 1j and 1n). This leads to highly anisotropic downstream ion distributions with $P_{i\perp}/P_{i\parallel}>8$, an artifact of the reduced dimensionality. Electron distributions exhibit a similar lack of parallel heating in $V_{ez}$ (Figure 1l and p), but rapid thermalization occurs in the perpendicular directions, $V_{ex}$ and $V_{ey}$, immediately downstream of the ramp (Figure 1c, 1k, 1o, and 1s).
As demonstrated later in the 2D quasi-perpendicular simulation (Figure 5), downstream parallel and oblique propagating waves, absent in the 1D case, play a crucial role in redistributing free energy associated with temperature anisotropy. A detailed discussion of these kinetic waves in the downstream region of quasi-perpendicular shocks is presented in Section 2.2.1.
@@
Figure 2: 1D simulation of a quasi-parallel shock with $\theta_{Bn}=30^{\circ}$ and $m_{i}/m_{e}=25$. The plasma moments and EM fields are normalized by the upstream quantities. Phase space distributions are normalized by the maximum phase space density in the displayed regions. (a-e) Close-up of plasma quantities and electromagnetic fields near the shock front at $t=19\,\omega_{ci,\text{up}}^{-1}$. (f) $x-t$ stack plot of the total magnetic field, overlaid with ion density contours. (g,h) Ion and electron $x-v_{x}$ phase space plot, with the $B_{z}$ profile shown in blue for reference. (i-l) Downstream ion and electron velocity space distributions. (m-p) Ion and electron velocity space distributions at the shock front. (q-t) Upstream ion and electron velocity space distributions.
In the 1D quasi-parallel shock run, we set a wider x extent from $-150\,d_{i,\text{up}}$ to $250\,d_{i,\text{up}}$ compared to the quasi-perpendicular case. This guarantees enough space for the evolution of reflected ions at the shock as well as the growth of upstream waves. Figure 2 illustrates the results of a 1D quasi-parallel shock simulation with $\theta_{Bn}=30^{\circ}$, presented in the same format as Figure 1. Notably, this simulation reveals significantly more dynamic wave activity than the quasi-perpendicular case, evident in both the plasma quantities (Figure 2a-c) and electromagnetic fields (Figure 2d-e). Within the upstream region ($x>-60\,d_{i,\text{up}}$), a dominant large-scale linearly polarized ultra-low frequency (ULF) wave extends far upstream and triggers perturbations in all variables. It possesses significant in-phase density and magnetic field perturbations (Figure 2a and 2d). Analysis of its phase velocity reveals that this wave mode propagate slightly slower than the inflow velocity in the lab frame. Transforming to the plasma rest frame (i.e., the frame moving with the inflow velocity), the phase velocities reverse direction, now pointing upstream in the +x direction. We thus identify this large-scale wave as the foreshock fast magnetosonic mode (e.g. J. Eastwood et al. [2005]). Corresponding to these wave fluctuations, Figure 2f, a spatiotemporal $x-t$ magnetic field stack plot, demonstrates the cyclic reformation of the main shock front from $t=10\,\omega_{ci}^{-1}$ to $20\,\omega_{ci}^{-1}$. This reformation is an unsteady process, driven by the nonlinear development of upstream waves. The match of the extension of upstream enhanced magnetic field with the density peaks, shown by the iso-density colored lines, again confirms its magnetosonic property. Furthermore, the downstream magnetic fluctuations exhibit higher amplitudes and broader spatial extents, consistent with observational data from Earth’s quasi-parallel magnetosheath (e.g., E. A. Lucek et al. [2002]).
@@
Within the upstream region ($x>-60\,d_{i,\text{up}}$), two distinct wave modes are clearly identified (Figure 2a, 2d, and 2e). The first mode consists of large-scale, linearly polarized waves with a wavelength of $\lambda\sim 40\,d_{i,\text{up}}\sim 5,300\,\text{km}$, a period of $T_{\text{w}}\sim 2\,(2\pi\omega_{ci}^{-1})\sim 26\,\text{s}$, and a wave amplitude of $\delta B/B_{0}\simeq 1$. These waves extend far upstream. The second mode comprises small-scale, left-hand polarized (in the lab frame) waves with a wavelength of $\lambda\lesssim 3.8\,d_{i,\text{up}}\sim 500\,\text{km}$, a period of $T_{\text{w}}\sim 1/4\,(2\pi\omega_{ci}^{-1})\sim 3\,\text{s}$, and a wave amplitude of $\delta B/B_{0}\lesssim 0.5$. These waves are localized on the leading edges of steepened structures. While both wave types exhibit magnetic field perturbations, only the large-scale mode induces significant in-phase density perturbations (Figure 2a and 2d). Analysis of their phase velocities reveals that both wave modes propagate slightly slower than the inflow velocity in the lab frame. Transforming to the plasma rest frame (i.e., the frame moving with the inflow velocity), the phase velocities reverse direction, now pointing upstream in the +x direction. Importantly, the left-hand polarization of the small-scale waves in the lab frame becomes right-hand polarized in the plasma rest frame. Based on these polarization and propagation characteristics, we identify the large-scale mode as the fast magnetosonic mode and the small-scale mode as the whistler mode, which in fact lies in the same branch in the cold plasma dispersion relation [S. P. Gary, 1993]. These magnetosonic-whistler waves are triggered by instabilities corresponding to the ion-ion beam interactions [S. P. Gary et al., 1984] but not driven by anisotropy.
@@
As the fast magnetosonic waves are convected back towards the shock by the super-Alfvénic upstream flow, they evolve into shocklets [M. Hoppe et al., 1981; L. B. Wilson III, 2016], i.e. steepened flanks characterized by small-scale fluctuations ($\delta B/B_{0}<2$) that function as miniature shocks. Short, Large Amplitude Magnetic Structures (SLAMS, or magnetic pulsations, S. J. Schwartz et al. [1992]) share similarities with shocklets but are defined by larger magnetic field fluctuations ($\delta B/B_{0}>2$) and potential soliton-like behavior. Consequently, our nonlinearly steepened ULF wave signatures align more closely with observed shocklets, where nonlinear evolution is driven by interactions with energetic particle pressure gradients, and the ensemble of these structures forms the shock transition (e.g., [N. Dubouloz & M. Scholer, 1995]). At this snapshot $t=19\,\omega_{ci,\text{up}}^{-1}$, a shocklet has just arrived at the main shock front. The normal velocity between $x=-65\,d_{i,\text{up}}$ and $-55\,d_{i,\text{up}}$ exhibits a step-like profile between upstream and downstream values, accompanied by a nearby density peak. Notably, whistler-like velocity and magnetic field fluctuations are identified in the foreshock region near the main shock front and trailing edges of density enhancements. The whistler critical Mach number depends on $m_{i}/m_{e}$ that affects the precursor existence condition. Larger mass ratios makes it easier to satisfy the criterion and trigger whistlers. Section 2.1.3 discusses on this point further.
It is crucial to acknowledge certain limitations inherent in the 1D planar shock setup. Firstly, with only one spatial degree of freedom, wave propagation is constrained to the x-direction. We established our background upstream magnetic field at a shock-normal angle $\theta_{Bn}=30^{\circ}$ relative to the -x direction. Consequently, this oblique wave angle remains fixed and may not align with the maximum wave growth direction predicted by linear theory. However, E. A. Lucek et al. [2002] reported from in-situ satellite measurements that the average $\theta_{Bn}$ is approximately $20^{\circ}$, a value inconsistent with the prediction for maximum parallel propagating linear wave growth. This discrepancy highlights the complexity of wave behavior in real-world environments compared to idealized simulations. Secondly, the periodic condition applied in the y and z direction limits the growth of structures that are intrinsically high-dimensional, such as downstream plasma jets and return flows. Further discussion with 2D results is presented in Section 2.2.2.
@@ 2.1.3 Impact of Mass Ratio
Figure 3: Influence of the ion-to-electron mass ratio $m_{i}/m_{e}$ on the normalized magnetic field profile across 1D collisionless shocks at $t=20\,\omega_{ci}^{-1}$. The grid resolution $\Delta x=1\,d_{e}$. Panels (a-c) show quasi-perpendicular shocks ($\theta_{Bn}=87^{\circ}$), while panels (d-f) show quasi-parallel shocks ($\theta_{Bn}=30^{\circ}$). The x-axes are shifted to align the shock fronts for easier comparison. The spatial extent shown is $50\,d_{i,\text{up}}$ for quasi-perpendicular shocks and $200\,d_{i,\text{up}}$ for quasi-parallel shocks.
[…]
@@ 2.1.4 Effect of Grid Resolution
Figure 4: Influence of the grid resolution $\Delta x$ on the normalized electromagnetic field profile across 1D collisionless shocks of $m_{i}/m_{e}=25$. Panels (a-h) show quasi-perpendicular shocks ($\theta_{Bn}=87^{\circ}$) at $t=20\,\omega_{ci}^{-1}$, while panels (i-p) show quasi-parallel shocks ($\theta_{Bn}=30^{\circ}$) at $t=20\,\omega_{ci}^{-1}$. The x-axes are shifted to align the shock fronts for easier comparison. The spatial extent shown is $50\,d_{i,\text{up}}$ for quasi-perpendicular shocks and $200\,d_{i,\text{up}}$ for quasi-parallel shocks.
To investigate the influence of grid resolution on the simulation results, we conducted a series of 1D simulations with a fixed mass ratio $m_{i}/m_{e}=25$ and varying grid resolutions $\Delta x$. Figure 4 compares the electromagnetic field profiles for fully developed quasi-perpendicular (a-d) and quasi-parallel (e-h) shocks, respectively, with $\Delta x=0.5,\,1,\,2,\,4\,d_{e,\text{up}}$. The coarsest resolution $\Delta x=4\,d_{e,\text{up}}\simeq 1\,d_{i,\text{up}}$ is close to the ion inertial length. Explicit full-PIC methods generally cannot work at this resolution due to severe numerical heating and instabilities. In the FLEKS quasi-perpendicular shock simulations, coarser resolutions lead to smoother EM field profiles, while the solutions stay meaningful at sub-ion scales. Similarly, for quasi-parallel shocks, coarser resolutions result in a weakening of the downstream and upstream field fluctuations, but large-scale structures persist. These observations highlight the asymptotic behavior in shock simulations across the sub-ion scales, demonstrating the stability of the semi-implicit PIC algorithm under different grid resolutions overcoming the constraints of explicit PIC methods.
@@ 2.2 Local 2D Simulations
To further validate the model, we extend the investigation to two dimensions (2D). These 2D simulations employ the same initial plasma parameters and proton-to-electron mass ratio of $m_{i}/m_{e}=25$ as the 1D cases, with the shock front initially normal to the x-direction and the magnetic field lying in the x-y plane. The simulation domain spans $-70\,d_{i,\text{up}}$ to $70\,d_{i,\text{up}}$ in the y direction. The x-extent is $200\,d_{i,\text{up}}\simeq 26,400\,\text{km}$ for the quasi-perpendicular shock case and $400\,d_{i,\text{up}}\simeq 52,800\,\text{km}$ for the quasi-parallel shock case, consistent with the 1D runs.
@@
Figure 5: 2D $\theta_{Bn}=87^{\circ}$ quasi-perpendicular shock run with $m_{i}/m_{e}=25$ at $t=20\,\omega_{ci,\text{up}}^{-1}$. The plasma moments and EM fields are normalized by the upstream quantities. Phase space distributions are normalized by the maximum phase space density in the displayed regions. (a-c) Zoomed-in view of magnetic field components near the shock front. (d-h) Plasma quantities and electromagnetic fields along the y=0 dashed line indicated in panel a. (i,j) Ion and electron $x-v_{x}$ phase space plots, with the $B_{y}$ profile shown in blue for reference. (k-n) Ion and electron velocity space distributions in the downstream region. (o-r) Ion and electron velocity space distributions at the shock front. (s-v) Ion and electron velocity space distributions ahead of the shock ramp. The y-extent for taking the velocity distributions goes from $-2\,d_{i,\text{up}}$ to $2\,d_{i,\text{up}}$.
@@
Remnants of the 1D downstream periodic structures are still discernible between $x=-50\,d_{i,\text{up}}$ and $-30\,d_{i,\text{up}}$ (Figure 5d, 5f, 5g), but their extent in the shock normal direction is limited to less than $20\,d_{i,\text{up}}$. The normalized magnetic fields in Figure 5a-c are presented on a symmetric logarithmic scale centered at 0. Immediately downstream of the shock front at $x=-30\,d_{i,\text{up}}$, strong fluctuations in the transverse $B_{x}$ and $B_{z}$ components are shown in the color contours, with amplitudes $\lesssim 2\,B_{\text{up}}$ and a wavelength of approximately $4\,d_{i,\text{up}}$. The comparable magnitudes of $B_{x}$ and $B_{z}$ oscillations suggest near-circular polarization, identifying these fluctuations as the parallel-propagating Alfvén ion-cyclotron (AIC) waves, also known as the electromagnetic ion cyclotron (EMIC) waves. These waves are generated by ion anisotropy shown in Figure 5o-p, and also contributes to the observed surface rippling along the y-direction.
@@
where $\beta_{i\parallel}=P_{i\parallel}/P_{B}$, $S$ and $\alpha$ are fitting parameters dependent on the wave growth rate $\gamma$ and plasma conditions. For simplicity, we adopt the fitting parameters from L. Blum et al. [2012] for $\gamma/\omega_{ci}=0.001$, yielding $S=0.429$ and $\alpha=0.409$. With $\beta_{\parallel}\simeq 2\mu_{0}P_{i,yy}/B_{\text{down}}^{2}\sim 1.9$, the EMIC instability criterion (3) is also marginally satisfied.
This linear instability analysis confirms that the conditions downstream of the 2D quasi-perpendicular shock are conducive to the generation of both mirror modes and EMIC waves. The 2D simulation, incorporating realistic solar wind conditions upstream of Earth [C. S. Salem et al., 2023], provides a more accurate representation of the plasma environment at Earth’s bow shock compared to the 1D case. This highlights the importance of at least 2D modeling for capturing the complex dynamics of quasi-perpendicular shocks, including the interplay of anisotropy-driven instabilities and wave generation.
@@
Figure 6: 2D $\theta_{Bn}=30^{\circ}$ quasi-parallel shock run with $m_{i}/m_{e}=25$ at $t=20\,\omega_{ci,\text{up}}^{-1}$. The plasma moments and EM fields are normalized by the upstream quantities. Phase space distributions are normalized by the maximum phase space density in the displayed regions. (a-c) Zoomed-in view of magnetic field components near the shock front. (d-h) Plasma quantities and electromagnetic fields along the $y=0$ dashed line indicated in panel a. (i-j) Ion and electron $x-V_{x}$ phase space density. (k-t) Ion velocity space distributions at five locations across the shock front, each spanning $10\,d_{i,\text{up}}$ in the x-direction and sampled over $-5\,d_{i,\text{up}}\leq y\leq 5\,d_{i,\text{up}}$.
In the 2D quasi-parallel shock simulation, the upstream magnetic field is oriented at a $30^{\circ}$ angle to the +x direction, tilted towards the +y direction in-plane. Initially the out-of-plane $B_{z}$ is zero. Figure 6 presents a detailed view of the 2D quasi-parallel shock results, following a similar format to Figure 5. The main shock front, located near $x=-80\,d_{i,\text{up}}$ at $t=20\,\omega_{ci,\text{up}}^{-1}$, exhibits surface waves along the tangential direction, distorting the initially planar shock surface. Unlike the 1D simulations, variations in $B_{x}$ are now present (Figure 6a). The upstream wave fronts are also tilted due to shock rippling (see Figure 9). Similar to the 1D case, the downstream region exhibits stronger, more irregular magnetic field fluctuations than the upstream region (Figure 6b-c). The dominant wave mode observed upstream is the large-scale fast magnetosonic wave, with wavelength $\lambda\sim 50\,d_{i,\text{up}}\sim 6,600\,\text{km}$, period $T_{\text{w}}\sim 1.5\,(2\pi\omega_{ci}^{-1})\sim 20\,\text{s}$ and $\delta B/B_{0}\simeq 1$, where $\delta B$ is the fluctuating magnetic field and $B_{0}$ is the background average magnetic field. Whistler-like velocity and magnetic field perturbations take place at the main shock front and the leading edges of steepened structures.

### v2

@@ 3.1 Local 1D Simulations
Three locations, centered downstream at $x=-2.5\,d_{i,\text{up}}$ (Figure 1i-l), across the shock at $x=2.5\,d_{i,\text{up}}$ (Figure 1m-p), and upstream in front of the shock at $x=7.5\,d_{i,\text{up}}$ (Figure 1q-t), each spanning $5\,d_{i,\text{up}}$, are selected to examine the velocity space distributions. The three-dimensional velocity space distributions are projected onto the $V_{x}-V_{y}$ and $V_{x}-V_{z}$ planes, where the z-axis is approximately aligned with the background magnetic field. The 1D simulation neglects parallel heating along the z-direction. Consequently, no proton thermalization occurs along $V_{iz}$ in the downstream and ramp regions (Figure 1j and 1n). This leads to highly anisotropic downstream ion distributions with $P_{i\perp}/P_{i\parallel}>8$. Electron distributions exhibit a similar lack of parallel heating in $V_{ez}$ (Figure 1l and p), but rapid thermalization occurs in the perpendicular directions, $V_{ex}$ and $V_{ey}$, immediately downstream of the ramp (Figure 1c, 1k, 1o, and 1s).
@@
Figure 2: 1D simulation of a quasi-parallel shock with $\theta_{Bn}=30^{\circ}$ and $m_{i}/m_{e}=25$. The plasma moments and EM fields are normalized by the upstream quantities. Phase space distributions are normalized by the maximum phase space density in the displayed regions. (a-e) Close-up of plasma quantities and electromagnetic fields near the shock front at $t=19\,\omega_{ci,\text{up}}^{-1}$. (f) xtx-t stack plot of the total magnetic field, overlaid with ion density contours. (g,h) Ion and electron $x-v_{x}$ phase space plot, with the $B_{z}$ profile shown in blue for reference. (i-l) Downstream ion and electron velocity space distributions. (m-p) Ion and electron velocity space distributions at the shock front. (q-t) Upstream ion and electron velocity space distributions.
In the 1D quasi-parallel shock run, we set a wider x extent from $150\,d_{i,\text{up}}$ to $250\,d_{i,\text{up}}$ compared to the quasi-perpendicular case. This guarantees enough space for the evolution of reflected ions at the shock as well as the growth of upstream waves. Figure 2 illustrates the results of a 1D quasi-parallel shock simulation with $\theta_{Bn}=30^{\circ}$, presented in the same format as Figure 1. Notably, this simulation reveals significantly more dynamic wave activity than the quasi-perpendicular case, evident in both the plasma quantities (Figure 2a-c) and electromagnetic fields (Figure 2d-e). Within the upstream region ($x>-60\,d_{i,\text{up}}$), a dominant large-scale linearly polarized ultra-low frequency (ULF) wave extends far upstream and triggers perturbations in all variables. It possesses significant in-phase density and magnetic field perturbations (Figure 2a and 2d). Analysis of its phase velocity reveals that this wave mode propagate slightly slower than the inflow velocity in the lab frame. Transforming to the plasma rest frame (i.e., the frame moving with the inflow velocity), the phase velocities reverse direction, now pointing upstream in the +x direction. We thus identify this large-scale wave as the foreshock fast magnetosonic mode (e.g. J. Eastwood et al. [2005]). Corresponding to these wave fluctuations, Figure 2f, a spatiotemporal xtx-t magnetic field stack plot, demonstrates the cyclic reformation of the main shock front from $t=10\,\omega_{ci,\text{up}}^{-1}$ to $20\,\omega_{ci,\text{up}}^{-1}$. This reformation is an unsteady process, driven by the nonlinear development of upstream waves. The match of the extension of upstream enhanced magnetic field with the density peaks, shown by the iso-density colored lines, again confirms its magnetosonic property. Furthermore, the downstream magnetic fluctuations exhibit higher amplitudes and broader spatial extents, consistent with observational data from Earth’s quasi-parallel magnetosheath (e.g., E. A. Lucek et al. [2002]).
@@
Within the upstream region ($x>-60\,d_{i,\text{up}}$), two distinct wave modes are clearly identified (Figure 2a, 2d, and 2e). The first mode consists of large-scale, linearly polarized waves with a wavelength of $\lambda\sim 40\,d_{i,\text{up}}\sim 5,300\,\text{km}$, a period of $T_{\text{w}}\sim 2\,(2\pi\omega_{ci,\text{up}}^{-1})\sim 26\,\text{s}$, and a wave amplitude of $\delta B/B_{0}\simeq 1$. These waves extend far upstream. The second mode comprises small-scale, left-hand polarized (in the lab frame) waves with a wavelength of $\lambda\lesssim 3.8\,d_{i,\text{up}}\sim 500\,\text{km}$, a period of $T_{\text{w}}\sim 1/4\,(2\pi\omega_{ci,\text{up}}^{-1})\sim 3\,\text{s}$, and a wave amplitude of $\delta B/B_{0}\lesssim 0.5$. These waves are localized on the leading edges of steepened structures. While both wave types exhibit magnetic field perturbations, only the large-scale mode induces significant in-phase density perturbations (Figure 2a and 2d). Analysis of their phase velocities reveals that both wave modes propagate slightly slower than the inflow velocity in the lab frame. Transforming to the plasma rest frame (i.e., the frame moving with the inflow velocity), the phase velocities reverse direction, now pointing upstream in the +x direction. Importantly, the left-hand polarization of the small-scale waves in the lab frame becomes right-hand polarized in the plasma rest frame. Based on these polarization and propagation characteristics, we identify the large-scale mode as the fast magnetosonic mode and the small-scale mode as the whistler mode, which in fact lies in the same branch in the cold plasma dispersion relation [S. P. Gary, 1993]. These magnetosonic-whistler waves are triggered by instabilities corresponding to the ion-ion beam interactions [S. P. Gary et al., 1984].
@@
At this snapshot $t=19\,\omega_{ci,\text{up}}^{-1}$, a shocklet has just arrived at the main shock front. The normal velocity between $x=-65\,d_{i,\text{up}}$ and $55\,d_{i,\text{up}}$ exhibits a step-like profile between upstream and downstream values, accompanied by a nearby density peak. Notably, whistler-like velocity and magnetic field fluctuations are identified in the foreshock region near the main shock front and trailing edges of density enhancements. The whistler critical Mach number depends on $m_{i}/m_{e}$ that affects the precursor existence condition. Larger mass ratios makes it easier to satisfy the criterion and trigger whistlers. Section 3.1.3 discusses on this point further.
@@ 3.1.3 Impact of Mass Ratio
Figure 3: Influence of the ion-to-electron mass ratio $m_{i}/m_{e}$ on the normalized magnetic field profile across 1D collisionless shocks at $t=20\,\omega_{ci,\text{up}}^{-1}$. The grid resolution $\Delta x=1\,d_{e}$. Panels (a-c) show quasi-perpendicular shocks ($\theta_{Bn}=87^{\circ}$), while panels (d-f) show quasi-parallel shocks ($\theta_{Bn}=30^{\circ}$). The x-axes are shifted to align the shock fronts for easier comparison. The spatial extent shown is $50\,d_{i,\text{up}}$ for quasi-perpendicular shocks and $200\,d_{i,\text{up}}$ for quasi-parallel shocks.
To assess the sensitivity of 1D shock structures to the ion-to-electron mass ratio $m_{i}/m_{e}$, we conduct simulations with a fixed grid resolution relative to the electron skin depth $d_{e}$, specifically $\Delta x=d_{e}$. We keep the real ion mass while setting different electron masses in the simulations. Figure 3 compares the magnetic field profiles for quasi-perpendicular (a-c) and quasi-parallel (d-f) shocks with $m_{i}/m_{e}=25,\,100,\,400$. Because the grid resolution is fixed with respect to $d_{e}$, the effective resolution in terms of the ion inertial length $d_{i}$ varies with the mass ratio as $\Delta x=d_{e}=d_{i}\sqrt{m_{e}/m_{i}}$, resulting in a corresponding change in the number of cells across the shock transition. Despite this difference in effective $d_{i}$ resolution, the overall magnetic field profiles remain remarkably similar across the different mass ratios, in both field perturbation magnitudes and periods.
@@ 3.1.4 Effect of Grid Resolution
Figure 4: Influence of the grid resolution $\Delta x$ on the normalized electromagnetic field profile across 1D collisionless shocks of $m_{i}/m_{e}=25$. Panels (a-h) show quasi-perpendicular shocks ($\theta_{Bn}=87^{\circ}$) at $t=20\,\omega_{ci,\text{up}}^{-1}$, while panels (i-p) show quasi-parallel shocks ($\theta_{Bn}=30^{\circ}$) at $t=20\,\omega_{ci,\text{up}}^{-1}$. The x-axes are shifted to align the shock fronts for easier comparison. The spatial extent shown is $50\,d_{i,\text{up}}$ for quasi-perpendicular shocks and $200\,d_{i,\text{up}}$ for quasi-parallel shocks.
To investigate the influence of grid resolution on the simulation results, we conducted a series of 1D simulations with a fixed mass ratio $m_{i}/m_{e}=25$ and varying grid resolutions $\Delta x$. Figure 4 compares the electromagnetic field profiles for fully developed quasi-perpendicular (a-d) and quasi-parallel (e-h) shocks, respectively, with $\Delta x=0.5,\,1,\,2,\,4\,d_{e,\text{up}}$. The coarsest resolution $\Delta x=4\,d_{e,\text{up}}\simeq 1\,d_{i,\text{up}}$ is close to the ion inertial length. In the quasi-perpendicular shock simulations, coarser resolutions lead to smoother EM field profiles, while the solutions stay meaningful at sub-ion scales. Similarly, for quasi-parallel shocks, coarser resolutions result in a weakening of the downstream and upstream field fluctuations, but large-scale structures persist. These observations demonstrate the asymptotic behavior in FLEKS shock simulations across the sub-ion scales.
@@ 3.2 Local 2D Simulations
To further validate the model, we extend the investigation to two dimensions (2D). These 2D simulations employ the same initial plasma parameters and proton-to-electron mass ratio of $m_{i}/m_{e}=25$ as the 1D cases, with the shock front initially normal to the x-direction and the magnetic field lying in the x-y plane. The simulation domain spans $70\,d_{i,\text{up}}$ to $70\,d_{i,\text{up}}$ in the y direction. The x-extent is $200\,d_{i,\text{up}}\simeq 26,400\,\text{km}$ for the quasi-perpendicular shock case and $400\,d_{i,\text{up}}\simeq 52,800\,\text{km}$ for the quasi-parallel shock case, consistent with the 1D runs.
@@
Figure 5: 2D $\theta_{Bn}=87^{\circ}$ quasi-perpendicular shock run with $m_{i}/m_{e}=25$ at $t=20\,\omega_{ci,\text{up}}^{-1}$. The plasma moments and EM fields are normalized by the upstream quantities. Phase space distributions are normalized by the maximum phase space density in the displayed regions. (a-c) Zoomed-in view of magnetic field components near the shock front. (d-h) Plasma quantities and electromagnetic fields along the y=0 dashed line indicated in panel a. (i,j) Ion and electron $x-v_{x}$ phase space plots, with the $B_{y}$ profile shown in blue for reference. (k-n) Ion and electron velocity space distributions in the downstream region. (o-r) Ion and electron velocity space distributions at the shock front. (s-v) Ion and electron velocity space distributions ahead of the shock ramp. The y-extent for taking the velocity distributions goes from $2\,d_{i,\text{up}}$ to $2\,d_{i,\text{up}}$.
@@
Remnants of the 1D downstream periodic structures are still discernible between $x=-50\,d_{i,\text{up}}$ and $30\,d_{i,\text{up}}$ (Figure 5d, 5f, 5g), but their extent in the shock normal direction is limited to less than $20\,d_{i,\text{up}}$. The normalized magnetic fields in Figure 5a-c are presented on a symmetric logarithmic scale centered at 0. Immediately downstream of the shock front at $x=-30\,d_{i,\text{up}}$, strong fluctuations in the transverse $B_{x}$ and $B_{z}$ components are shown in the color contours, with amplitudes $\lesssim 2\,B_{\text{up}}$ and a wavelength of approximately $4\,d_{i,\text{up}}$. The comparable magnitudes of $B_{x}$ and $B_{z}$ oscillations suggest near-circular polarization, identifying these fluctuations as the parallel-propagating Alfvén ion-cyclotron (AIC) waves, also known as the electromagnetic ion cyclotron (EMIC) waves. These waves are generated by ion anisotropy shown in Figure 5o-p, and also contributes to the observed surface rippling along the y-direction.
@@
where $\beta_{i\parallel}=P_{i\parallel}/P_{B}$, $S$ and $\alpha$ are fitting parameters dependent on the wave growth rate $\gamma$ and plasma conditions. For simplicity, we adopt the fitting parameters from L. Blum et al. [2012] for $\gamma/\omega_{ci,\text{up}}=0.001$, yielding $S=0.429$ and $\alpha=0.409$. With $\beta_{\parallel}\simeq 2\mu_{0}P_{i,yy}/B_{\text{down}}^{2}\sim 1.9$, the EMIC instability criterion (3) is also marginally satisfied. This linear instability analysis confirms that the conditions downstream of the 2D quasi-perpendicular shock are conducive to the generation of both mirror modes and EMIC waves.
@@
Figure 6: 2D $\theta_{Bn}=30^{\circ}$ quasi-parallel shock run with $m_{i}/m_{e}=25$ at $t=20\,\omega_{ci,\text{up}}^{-1}$. The plasma moments and EM fields are normalized by the upstream quantities. Phase space distributions are normalized by the maximum phase space density in the displayed regions. (a-c) Zoomed-in view of magnetic field components near the shock front. (d-h) Plasma quantities and electromagnetic fields along the $y=0$ dashed line indicated in panel a. (i-j) Ion and electron $x-V_{x}$ phase space density. (k-t) Ion velocity space distributions at five locations across the shock front, each spanning $10\,d_{i,\text{up}}$ in the x-direction and sampled over $5\,d_{i,\text{up}}\leq y\leq 5\,d_{i,\text{up}}$.
In the 2D quasi-parallel shock simulation, the upstream magnetic field is oriented at a $30^{\circ}$ angle to the +x direction, tilted towards the +y direction in-plane. Initially the out-of-plane $B_{z}$ is zero. Figure 6 presents a detailed view of the 2D quasi-parallel shock results, following a similar format to Figure 5. The main shock front, located near $x=-80\,d_{i,\text{up}}$ at $t=20\,\omega_{ci,\text{up}}^{-1}$, exhibits surface waves along the tangential direction, distorting the initially planar shock surface. Unlike the 1D simulations, variations in $B_{x}$ are now present (Figure 6a). The upstream wave fronts are also tilted due to shock rippling (see Figure 9). Similar to the 1D case, the downstream region exhibits stronger, more irregular magnetic field fluctuations than the upstream region (Figure 6b-c). The dominant wave mode observed upstream is the large-scale fast magnetosonic wave, with wavelength $\lambda\sim 50\,d_{i,\text{up}}\sim 6,600\,\text{km}$, period $T_{\text{w}}\sim 1.5\,(2\pi\omega_{ci,\text{up}}^{-1})\sim 20\,\text{s}$ and $\delta B/B_{0}\simeq 1$, where $\delta B$ is the fluctuating magnetic field and $B_{0}$ is the background average magnetic field. Whistler-like velocity and magnetic field perturbations take place at the main shock front and the leading edges of steepened structures.